hamidiso23:

Solve the following equation for aa. Be sure to take into account whether a letter is capitalized or not. D= Ga + 3h3a a= ??

hamidiso23:

btw the 3 after the h is an exponent

hamidiso23:

its due in 5 mins btw

SmokeyBrown:

Just want to clarify, the question says to solve for "aa". Is this supposed to mean "a", or are we meant to solve for "aa", because I don't see such a term in the equation (I guess it could mean a multiplied by a?) Also, I'm not sure what the significance of the capital letters D and G are; do you have any context for these variables by any chance? In any case, once we know for sure what we're solving for, we should be able to do so using algebra. Just so everything is clear and unambiguous, the equation we're working with looks like this: \[D = G*a + 3*h^3*a\] With multiplication between G and a, and multiplication between 3, h^3, and a, correct?

SmokeyBrown:

@hamidiso23 wrote:
btw the 3 after the h is an exponent
I'll be hasty, since I know you're on a time limit. By the way, this isn't part of a quiz or test, right?

hamidiso23:

@smokeybrown wrote:
@hamidiso23 wrote:
btw the 3 after the h is an exponent
I'll be hasty, since I know you're on a time limit. By the way, this isn't part of a quiz or test, right?
hw

hamidiso23:

@smokeybrown wrote:
Just want to clarify, the question says to solve for "aa". Is this supposed to mean "a", or are we meant to solve for "aa", because I don't see such a term in the equation (I guess it could mean a multiplied by a?) Also, I'm not sure what the significance of the capital letters D and G are; do you have any context for these variables by any chance? In any case, once we know for sure what we're solving for, we should be able to do so using algebra. Just so everything is clear and unambiguous, the equation we're working with looks like this: \[D = G*a + 3*h^3*a\] With multiplication between G and a, and multiplication between 3, h^3, and a, correct?
correct

SmokeyBrown:

For now, I'll assume that everything is as it seems We can simplify the equation: D = a*(G + 3h^3) Then, we can divide both sides of the equation by (G + 3h^3) to get: a = D/(G + 3h^3) And there you've technically solved for "a", which I assume is the kind of response you're expected to give for this question

hamidiso23:

@smokeybrown wrote:
For now, I'll assume that everything is as it seems We can simplify the equation: D = a*(G + 3h^3) Then, we can divide both sides of the equation by (G + 3h^3) to get: a = D/(G + 3h^3) And there you've technically solved for "a", which I assume is the kind of response you're expected to give for this question
yh

hamidiso23:

whats a =

SmokeyBrown:

We can find that the right side of the equation, (G∗a+3∗h3∗a), is equal to a*(G + 3h^3) (To show this, you can try distributing the terms over the parentheses with multiplication) Giving us the full equation D = a*(G + 3h^3) Divide both sides by (G + 3h^3) to get the equation: a = D/(G + 3h^3) That is what a equals. The actual numerical value of a depends on the values of D, G, and h, but that is what "a" is equal in terms of the other variables

hamidiso23:

a=D/G3h

SmokeyBrown:

As I said, the result I got from the algebra I showed is a = D/(G + 3h^3) You can see the steps I took to get this result, and I'll be happy to explain any of the details further if you'd like

hamidiso23:

tysm

SmokeyBrown:

Not a problem, glad I could help :)